1) Candiland is a closed economy, that does not trade with the rest of the world. Their autonomous consumption expenditure is $80 million, and the marginal propensity to consume is 0.5. Investment spending is constant at $100 million, and government expenditure is constant at $80 million. There are no income taxes.

a) What is the autonomous aggregate expenditure i.e. AE0?
b) What is the slope of the aggregate expenditure i.e. AE1?
c) What is the size of the multiplier in this economy?
d) What is the value of aggregate planned expenditure when real GDP is $800 million?
e) What can you say about the inventories when real GDP is $800 million?
f) What is the economy's equilibrium aggregate expenditure?
g) If a stock market boom increases autonomous consumption expenditure to $100 million, how would the equilibrium aggregate expenditure change?
1
Expert's answer
2019-08-07T09:15:19-0400

a) In closed economy AE= C+I+G

C= $80million + 0.5(Y). AE= $80 million + 0.5Y + $100 million + $80 million. aggregate expenditure = AE= $260 million + 0.5Y

b) Y= 0, AE= $260 million + 0.5(0) → autonomous expenditure is $260 million

c) the multiplier = "\frac{1}{1-0.5}" = 2


d) When real GDP is $800 million, AE= $260 million + 0.5 × $800 million, which is $660 million.

e) inventories will increase.

f) Y = $260million + 0.5Y = 0.5Y = $260 million. Y = "\frac{260}{1-0.5}" = $520 million.

g) AE= $100 million + 0.5Y + $100 million + $80 million= $280+0.5Y = $560. The equilibrium aggregate expenditure will increase.